# SequencePAR: Understanding Pedestrian Attributes via A Sequence Generation Paradigm

### v1

@@ III-C Network Architecture
Attribute Encoding Network.  To help the models better understand the pedestrian attributes that need to be classified, in this work, we take the raw attribute phases $\mathcal{A}_{set}=\{\mathcal{A}_{1},\mathcal{A}_{2},...,\mathcal{A}_{M}\}$ as the input. To be specific, we expand and split each attribute phase $\mathcal{A}_{i}$ into text representations with the help of prompt engineering. For example, the attribute "age $\leq$ 40" is firstly processed to "age less than 40", then, transformed into a sentence "the age of this pedestrian is less than 40 years old". After that, we obtain the text tokens $\mathcal{E}_{set}=\{\mathcal{E}_{1},\mathcal{E}_{2},...,\mathcal{E}_{M}\}$ using word embedding. Then, we adopt the text encoder of CLIP model [24] to learn the high-level semantic features of human attributes. Therefore, we can get the attribute query tokens $\{\mathcal{Q}_{1},\mathcal{Q}_{2},...,\mathcal{Q}_{M}\}$, which learns the concept of each attribute well. We believe this will make our PAR model better understand these attributes and predict the human attributes in a more accurate way.
Sequence Generation Network.  In this work, we formulate the pedestrian attribute recognition task as a novel sequence generation problem. Given the attribute query tokens $\mathcal{Q}_{i},i\in\{1,2,...,M\}$, we first introduce the position encoding tokens and add these tokens as the input of the sequence generation network. Specifically, the query tokens are treated as Key (K), Value (V), and Query (Q) and feed into the masked multi-head attention module. It receives the predicted attributes from the preceding block and adopts the casual mask to ensure that the predicted pedestrian attributes only depend on its previous sequence cues. As shown in Fig. 2, the output attribute $\mathcal{A}_{i}$ only attends to the attributes less than $i$, with the guidance of the attention mask. Then, the normalization layers and skip connections are adopted to process the obtained features and the output will be treated as the query input of the subsequent multi-head self-attention layer. The visual tokens are used as its key and value inputs. Then, the normalization and feed-forward layers are utilized to further enhance the decoded features $U$.
@@
IV Experiments
In this section, we will first introduce the datasets and evaluation metric in subsection IV-A. The implementation details are given in subsection IV-B. After that, we report our recognition results and compare them with other state-of-the-art algorithms in subsection IV-C. Then, we conduct extensive studies on our newly proposed SequencePAR framework in subsection IV-D. The visualization is also provided in subsection IV-E to help the readers better understand our model. Then, we discuss the difference between our model and existing recurrent neural network-based PAR models and also the limitation analysis in subsection IV-F and IV-G, respectively.
IV-A Datasets and Evaluation Metric
@@
-- PETA dataset [50] contains 19,000 outdoor or indoor pedestrian images and 61 binary attributes. These images are split into training, validation, and testing subset, which contains 9500, 1900, and 7600 images, respectively. In our experiments, we select 35 pedestrian attributes by following the work [50].
@@
-- PA100K dataset [29] is the largest pedestrian attribute recognition dataset which contains 100,000 pedestrian images, and 26 binary attributes. In our experiments, we split them into a training and validation set which contains 90,000 images, and a testing subset with the rest 10,000 images.
@@
-- RAP-ZS dataset is developed based on RAPv2 dataset and contains 17062, 4628, and 4928 pedestrian images for training, validation, and testing, respectively. No shared personal identity between the training and inference data. In our experiments, we select 53 attributes for the evaluation by following Jia et al. [53].
@@
where TP is predicting the correct positive sample, TN is predicting the correct negative sample, FP is a negative sample of prediction errors, and FN is a positive sample of prediction errors. The formulation of Precision, Recall, and F1 measure can be expressed as:
@@
TABLE I: Comparison with state-of-the-art methods on PETA and PA100K datasets. The first and second are shown in red and blue, respectively. "-" means this indicator is not available.
@@
MsVAA [56] ECCV 2018 ResNet101 78.56 86.79 86.12 86.46 - - - -
RA [57] AAAI 2019 Inception-V3 - 84.69 88.51 86.56 - - - -
VRKD [58] IJCAI 2019 ResNet50 80.95 88.37 87.47 87.91 78.49 88.42 86.08 87.24
AAP [59] IJCAI 2019 ResNet50 79.95 87.58 87.73 87.65 78.30 89.49 84.36 86.85
@@
IAA-Caps [63] PR 2022 OSNet 78.04 86.08 85.80 85.64 80.31 88.36 88.01 87.80
MCFL [64] NCA 2022 ResNet50 78.89 84.57 88.84 86.65 77.80 85.11 88.20 86.62
DRFormer [22] NC 2022 ViT-B/16 81.30 85.68 91.08 88.30 80.27 87.60 88.49 88.04
@@
CAS-SAL-FR [67] IJCV 2022 ResNet50 79.93 87.03 87.33 87.18 79.64 86.81 87.79 85.18
VTB [18] TCSVT 2022 ViT-B/16 79.60 86.76 87.17 86.71 80.89 87.88 89.30 88.21
FEMDAR [68] SPIC 2023 ResNet50 78.45 86.79 85.69 85.90 79.65 87.99 87.45 87.32
EALC [69] NC 2023 EfficientNet-B4 81.71 88.58 88.23 88.40 80.27 87.32 88.98 88.14
APTM [8] ACM MM 2023 Swin-B - - - - 80.17 88.31 87.84 88.07
@@
VTB* [18] TCSVT 2022 ViT-L/14 79.59 86.66 87.82 86.97 81.76 87.87 90.67 88.86
SequencePAR - ViT-L/14 84.92 90.44 90.73 90.46 83.94 90.38 90.23 90.10
TABLE II: Comparison with state-of-the-art methods on RAPv1 and RAPv2 datasets. The first and second are shown in red and blue, respectively. "-" means this indicator is not available.
@@
MsVAA [56] ECCV 2018 ResNet101 - - - - 65.57 77.37 79.17 78.26
RA [57] AAAI 2019 Inception-V3 - 79.45 79.23 79.34 - - - -
VRKD [58] IJCAI 2019 ResNet50 69.79 82.13 80.35 81.23 - - - -
AAP [59] IJCAI 2019 ResNet50 68.37 81.04 80.27 80.65 - - - -
@@
ALM [61] ICCV 2019 BN-Inception 68.17 74.71 86.48 80.16 64.79 73.93 82.03 77.77
JLAC [14] AAAI 2020 ResNet50 69.15 79.31 82.40 80.82 64.42 75.69 79.18 77.40
SSCsoft [62] ICCV 2021 ResNet50 68.37 75.05 87.49 80.43 - - - -
IAA-Caps [63] PR 2022 OSNet 68.47 79.56 82.06 80.37 - - - -
MCFL [64] NCA 2022 ResNet50 67.28 73.44 87.75 79.96 - - - -
DRFormer [22] NC 2022 ViT-B/16 70.60 80.12 82.77 81.42 - - - -
@@
CAS-SAL-FR [67] IJCV 2022 ResNet50 68.59 77.56 83.81 80.56 - - - -
VTB [18] TCSVT 2022 ViT-B/16 69.44 78.28 84.39 80.84 67.48 76.41 83.32 79.35
FEMDAR [68] SPIC 2023 ResNet50 66.88 79.11 79.24 78.76 - - - -
EALC [69] NC 2023 EfficientNet-B4 69.65 79.82 83.61 81.67 - - - -
@@
VTB* [18] TCSVT 2022 ViT-L/14 69.78 78.09 85.21 81.10 67.58 76.19 84.00 79.52
SequencePAR - ViT-L/14 71.47 82.40 82.09 82.05 70.14 81.37 81.22 81.10
TABLE III: Comparison with state-of-the-art methods on PETA-ZS and RAP-ZS datasets.
@@
MCFL[64] NCA 2022 ResNet50 57.04 68.47 74.35 71.29 63.37 71.21 83.86 77.02
VTB [18] TCSVT 2022 ViT-B/16 60.50 73.29 74.40 73.38 64.73 74.93 80.85 77.35
VTB*[18] TCSVT 2022 ViT-L/14 63.12 74.77 77.24 75.50 68.34 76.81 84.51 80.07
SequencePAR - ViT-L/14 66.70 78.75 78.52 78.40 70.28 82.13 80.55 81.14
IV-B Implementation Details
@@
In our experiments, we adopt the ViT-L/14 version of the CLIP [24] model as our feature extractor. The vision encoder of our model is a 24-layer vision Transformer [17] network and the dimension of the hidden layer is 1024. Our sequence generation network is a 6-layer Transformer Decoder, each with eight attention headers. The dimension of the hidden layer is 768. Considering the pedestrian images are all long strip shapes, however, the resolution of input images for the CLIP is $224\times 224$. To adapt to the input resolution, we padded the raw pedestrian images using black pixels into a square shape and resized them into $224\times 224$. Note that, the randomly crop and flip operations are also adopted for data augmentation.
@@
During the training phase, we utilize the ground truth as the input and predict the attributes in a parallel manner, which can reduce the cumulative error significantly and make our model converge faster. For the model inference, we adopt a step-by-step generation approach that uses the output predicted in the previous step and historical output as the inputs of our network. For the detailed parameters, we set the learning rate as 1e-5. We train our model for 50 epochs with the Adam optimizer [72] and set the batch size as 32. Our source code is implemented using Python and the deep learning framework PyTorch [73]. The experiments are conducted based on a server with GPU A100. More details can be found in our source code.
IV-C Comparison with Other SOTA Models
@@
In this section, we will report our recognition results on all six datasets and compare them with existing state-of-the-art pedestrian attribute recognition algorithms. Note that the results of VTB* are obtained by replacing the backbone network of VTB using the visual encoder of CLIP model.
@@
Results on PETA dataset.  As shown in Table I, our proposed SequencePAR model achieves the best performance on most of the evaluation metrics, i.e., 84.92, 90.44, 90.73, 90.46 on the Accuracy, Precision, Recall, and F1-measure, respectively. Compared with other Transformer-based pedestrian attribute recognition models, we can find that our model exceeds both the VTB (ViT-base/16, 79.60/86.76/87.17/86.71) and VTB* (ViT-L/14, 79.59/86.66/87.82/86.97) by +5.32/+3.68/+3.56/+3.67 and +5.33/+3.78/+2.91/+3.41. Our model also achieves better performance than the DRFormer [22] on the Accuracy, Precision, and F1-measure metric. Therefore, we can draw the conclusion that our model achieves state-of-the-art results on the PETA dataset, which fully validates the superiority of our sequence generation framework for the pedestrian attribute recognition task.
@@
Results on PA100K dataset.  As shown in Table I, we can find that our model SequencePAR achieves 83.94/90.38/90.23/90.10 on the Accuracy/Precision/Recall/F1 metric, respectively. These results are better than the multi-label classification model VTB* (81.76/87.87/90.67/88.86) which also adopts the ViT-L/14 model as the backbone network. These experiments demonstrate that our attribute generation-based scheme works better than the discriminative-based learning framework. Also, our model achieves better results than the compared state-of-the-art recognition models.
@@
Results on RAPv1 dataset.  As shown in Table II, we can find that our model achieves 71.47/82.40/82.09/82.02 on the Accuracy/Precision/Recall/F1 metric, respectively. In contrast, the strong baseline VTB* which is the discriminative method achieves 69.78/78.09/85.21/81.10 on these metrics. Obviously, our results are better than theirs, and also the compared attribute recognition models, including CAS-SAL-FR  [67] (IJCV-2022), DRFormer [22] (NC 2022). These experiments fully validated the effectiveness of our proposed model for pedestrian attribute recognition.
@@
Results on RAPv2 dataset.  As shown in Table II, we can find that the multi-label classification based attribute recognition model VTB* (backbone network ViT-L/14) achieves 67.58/76.19/84.00/79.52 on the Accuracy/Precision/Recall/F1 metric, respectively. Meanwhile, our model achieves 70.14, 81.37, 81.22, 81.10. Therefore, we can draw the conclusion that our model is better than the strong discriminative learning based pedestrian attribute recognition algorithm.
@@
Results on PETA-ZS dataset.  In addition to the aforementioned standard evaluation, in this work, we also conduct experiments on the zero-shot setting based datasets. According to the experimental results reported in Table III, we can find that our model achieves 66.70/78.75/78.52/78.40 on the Accuracy/Precision/Recall/F1 metric, respectively. These results are better than the multi-label classification model VTB* (63.12/74.77/77.24/75.50) which also adopts the ViT-L/14 model as the backbone network.
@@
Results on RAP-ZS dataset.  For the experimental results reported on the RAP-ZS dataset as illustrated in Table III, we can find that the baseline method VTB* which achieves 68.34/76.81/84.51/80.07 on the Accuracy/Precision/Recall/F1 metric, respectively. However, we get better results on these metrics, i.e., 70.28, 82.13, 80.55, 81.14. Based on the experiments on the PETA-ZS and RAP-ZS datasets, we can find that our model also works better in this setting. We believe this may be because we use a large visual-language model, which has better generalization ability. Another key reason is that the generative learning based strategy works better for pedestrian attribute recognition.
IV-D Ablation Study
Greedy Search or Beam Search for Attribute Generation? For the sequence generation tasks, like machine translation and image/video captioning, the greedy search and beam search are all widely used in the inference phase. The greedy search policy will select and remain the element with the maximum response score. In contrast, the beam search policy will always select and retain a fixed number of candidates (also called beam width $\mathcal{B}$) for each time step, and select the trajectory with the maximum summarised response.
In this part, we test different search policies to check whether they still work for pedestrian attribute generation. Specifically, we can find that the greedy search achieves $84.74,90.38,90.42,90.28$ on the Accuracy, Precision, Recall, and F1 measure, as shown in Table IV. When beam search is adopted, the recognition performance on the PETA dataset is similar to the greedy search, however, no improvements are obtained. Therefore, we adopt the simple greedy search scheme for other experiments in this work.
@@ IV-D Ablation Study
TABLE IV: Compare the greedy search and beam search with different beam widths. The best results are highlighted in bold.
@@
Results using Different Layers of Decoders. As shown in Table V, we also conduct experiments on the PETA dataset to check the influence of different layers used in our decoder network. To be specific, we set the layers as 1, 3, 6, 9, and 12. From Table V, we can find that the F1 measures are 89.80, 89.68, 90.46, 89.60, and 89.76, which are relatively stable. When 6 decoding layers are used, we get the best recognition performance on the PETA dataset. Therefore, we default set six layers as our decoder network when conducting other experiments in this paper.
TABLE V: Comparison of Different Decoding Layers for PAR on PETA dataset. The best results are highlighted in bold.

### v2

@@ 3.3 Network Architecture
Attribute Encoding Network.  To help the models better understand the pedestrian attributes that need to be classified, in this work, we take the raw attribute phases $\mathcal{A}_{set}=\{\mathcal{A}_{1},\mathcal{A}_{2},...,\mathcal{A}_{M}\}$ as the input. To be specific, we expand and split each attribute phase $\mathcal{A}_{i}$ into text representations with the help of prompt engineering. For example, the attribute ”age $\leq$ 40” is firstly processed to ”age less than 40”, then, transformed into a sentence ”the age of this pedestrian is less than 40 years old”. After that, we obtain the text tokens $\mathcal{E}_{set}=\{\mathcal{E}_{1},\mathcal{E}_{2},...,\mathcal{E}_{M}\}$ using word embedding. Then, we adopt the text encoder of CLIP model [24] to learn the high-level semantic features of human attributes. Therefore, we can get the attribute query tokens $\{\mathcal{Q}_{1},\mathcal{Q}_{2},...,\mathcal{Q}_{M}\}$, which learns the concept of each attribute well. Due to the strong mutual exclusivity or co-occurrence relationships among many attributes, for instance, in the PETA dataset, attributes such as casual top, formal top, casual trousers, formal trousers, and leather shoes exhibit significant interdependencies. These relationships are difficult to capture using visual features alone. Modeling such dependencies helps the system avoid inconsistent predictions, such as simultaneously predicting casual top and formal trousers. To address this, we encode attribute descriptions using the CLIP text encoder as queries, enabling the model to learn semantic dependencies among attributes through causal masking during training.
Sequence Generation Network.  In this work, we formulate the pedestrian attribute recognition task as a novel sequence generation problem. Given the attribute query tokens $\mathcal{Q}_{i},i\in\{1,2,...,M\}$, these represent text embeddings obtained by processing attribute words through prompt templates to generate complete sentences, which are then encoded by our text encoder. We use these embeddings as attribute query tokens within our framework. We first introduce the position encoding tokens and add these tokens as the input of the sequence generation network. Specifically, the query tokens are treated as Key (K), Value (V), and Query (Q) and feed into the masked multi-head attention module. This refers to our implementation of attention masking in the self-attention mechanism of the decoder. This method effectively prevents the model from accessing future information during autoregressive predictions. It receives the predicted attributes from the preceding block and adopts the causal mask to ensure that the predicted pedestrian attributes only depend on its previous sequence cues. As shown in Figure 2, the output attribute $\mathcal{A}_{i}$ only attends to the attributes less than $i$, with the guidance of the attention mask. Then, the normalization layers and skip connections are adapted to process the obtained features and the output will be treated as the query input of the subsequent multi-head self-attention layer. The visual tokens are used as the key and value inputs. Then, the normalization and feed-forward layers are utilized to further enhance the decoded features $U$.
@@
3.4 Training and Inference
As shown in Figure 3, we formulate pedestrian attribute recognition as a closed-set vocabulary sequence generation problem, where the vocabulary consists of all predefined attribute labels along with special tokens, such as the start-of-sequence <<BOS>> and end-of-sequence <<EOS>> markers. During training, the ground-truth attribute labels corresponding to each image are first converted into textual embeddings using a predefined prompt template. These embeddings are then right-shifted and fed into a Transformer decoder as input queries. Simultaneously, the visual features extracted by a CLIP encoder serve as conditional context for the decoder. The model is trained to predict the probability distribution of the next token in an autoregressive manner at each decoding step. During inference, the decoder receives the start token and the image’s visual context as input and generates a sequence of attribute indices step-by-step via greedy decoding. At each step, the model uses the previously generated index (converted into an embedding) together with the visual features to predict the next index. The resulting sequence of attribute indices is mapped into a one-hot vector representing the predicted attribute set. Special tokens are removed during post-processing, and duplicated indices are merged to form the final prediction.
Figure 3: The training and inference pipeline of the proposed SequencePAR framework.
4 Experiments
In this section, we will first introduce the datasets and evaluation metric in subsection 4.1. The implementation details are given in subsection 4.2. After that, we report our recognition results and compare them with other state-of-the-art algorithms in subsection 4.3. Then, we conduct extensive studies on our newly proposed SequencePAR framework in subsection 4.4. The visualization is also provided in subsection 4.5 to help the readers better understand our model. Then, we discuss the difference between our model and existing recurrent neural network-based PAR models and also the limitation analysis in subsection 4.7, respectively.
4.1 Datasets and Evaluation Metric
@@
- PETA dataset [5] contains 19,000 outdoor or indoor pedestrian images and 61 binary attributes. These images are split into training, validation, and testing subsets, which contains 9500, 1900, and 7600 images, respectively. In our experiments, we select 35 pedestrian attributes by following the work [5].
@@
- PA100K dataset [19] is the largest pedestrian attribute recognition dataset which contains 100,000 pedestrian images, and 26 binary attributes. In our experiments, we split them into a training and validation set which contains 90,000 images, and a testing subset with the remaining 10,000 images.
@@
- RAP-ZS dataset is constructed based on RAPv2 and consists of 17,062 images for training, 4,628 for validation, and 4,928 for testing. Importantly, there is no overlap in pedestrian identities between the training and inference sets, ensuring a strict zero-shot evaluation protocol. In our experiments, we follow Jia et al. [12] and select 53 attributes for evaluation.
@@
where TP is predicting the correct positive sample, TN is predicting the correct negative sample, FP is a negative sample of prediction errors, and FN is a positive sample of prediction errors. The formulation of Precision, Recall, and F1 measures can be expressed as:
@@
Table 1: Comparison with state-of-the-art methods on PETA and PA100K datasets. The first and second are shown in red and blue, respectively. ”-” means this indicator is not available.
@@
DRFormer [28] NC 2022 ViT-B//16 81.30 85.68 91.08 88.30 80.27 87.60 88.49 88.04
@@
VTB [2] TCSVT 2022 ViT-B//16 79.60 86.76 87.17 86.71 80.89 87.88 89.30 88.21
@@
VTB* [2] TCSVT 2022 ViT-L//14 79.59 86.66 87.82 86.97 81.76 87.87 90.67 88.86
SSPNet [26] PR 2024 Swin-S 82.80 88.48 90.55 89.50 80.63 87.79 89.32 88.55
ViT-RE++ [27] TMM 2024 ViT-B 81.64 88.59 88.82 88.70 81.47 89.78 89.77 88.88
SOFA [35] AAAI 2024 ViT-B 81.06 87.77 88.35 87.83 81.14 88.39 88.98 88.34
HDFL [33] NN 2025 ViT-B 79.66 87.08 87.16 86.85 80.23 87.45 88.74 87.72
AAR [32] NC 2025 ViT-B 82.46 89.04 89.01 88.81 81.96 88.78 89.62 88.87
SOFAFormer++ [34] TCSVT 2025 ViT-B 80.91 87.17 88.74 87.70 80.80 87.62 89.37 88.15
PIL [39] IJCV 2025 ConvNeXt-base - - - 87.49 - - - 89.53
SequencePAR - ViT-L//14 84.92 90.44 90.73 90.46 83.94 90.38 90.23 90.10
Table 2: Comparison with state-of-the-art methods on RAPv1 and RAPv2 datasets. The first and second are shown in red and blue, respectively. ”-” means this indicator is not available.
@@
DRFormer [28] NC 2022 ViT-B//16 70.60 80.12 82.77 81.42 - - - -
@@
VTB [2] TCSVT 2022 ViT-B//16 69.44 78.28 84.39 80.84 67.48 76.41 83.32 79.35
@@
VTB* [2] TCSVT 2022 ViT-L//14 69.78 78.09 85.21 81.10 67.58 76.19 84.00 79.52
SSPNet [26] PR 2024 Swin-S 70.21 80.14 82.90 81.50 - - - -
ViT-RE++ [27] TMM 2024 ViT-B 69.45 81.18 80.80 80.99 - - - -
SOFA [35] AAAI 2024 ViT-B 70.03 79.99 83.03 81.15 68.62 78.00 83.14 80.15
SOFAFormer++ [34] TCSVT 2025 ViT-B 69.73 78.77 84.05 80.97 67.86 77.57 83.81 79.66
HDFL [32] NN 2025 ViT-B 70.64 80.94 82.82 81.55 - - - -
AAR [33] NC 2025 ViT-B 70.49 80.25 83.55 81.51 68.22 78.30 82.18 79.85
PIL [39] IJCV 2025 ConvNeXt-base - - - 80.47 - - - -
SequencePAR - ViT-L//14 71.47 82.40 82.09 82.05 70.14 81.37 81.22 81.10
Table 3: Comparison with state-of-the-art methods on PETA-ZS and RAP-ZS datasets.
@@
VTB [2] TCSVT 2022 ViT-B//16 60.50 73.29 74.40 73.38 64.73 74.93 80.85 77.35
VTB*[2] TCSVT 2022 ViT-L//14 63.12 74.77 77.24 75.50 68.34 76.81 84.51 80.07
SOFA [35] AAAI 2024 ViT-B 62.07 74.97 75.13 74.63 66.26 78.20 79.44 78.42
SOFAFormer++ [34] TCSVT 2025 ViT-B 61.41 74.09 75.73 74.10 66.09 77.53 78.33 78.33
HDFL [32] NN 2025 ViT-B 62.01 75.07 75.36 74.78 66.70 78.52 79.81 78.42
AAR [33] NC 2025 ViT-B 62.89 75.93 75.66 75.38 66.51 78.83 79.13 78.60
PIL [39] IJCV 2025 ConvNeXt-base - - - 75.81 - - - 79.75
SequencePAR - ViT-L//14 66.70 78.75 78.52 78.40 70.28 82.13 80.55 81.14
4.2 Implementation Details
@@
In our experiments, we adopt the ViT-L//14 version of the CLIP [24] model as our feature extractor. The vision encoder of our model is a 24-layer vision Transformer [6] network and the dimension of the hidden layer is 1024. Our sequence generation network is a 6-layer Transformer Decoder, each with eight attention heads. The dimension of the hidden layer is 768. Considering the pedestrian images are all long strip shapes, however, the resolution of the input images for the CLIP is $224\times 224$. To adapt to the input resolution, we padded the raw pedestrian images using black pixels into a square shape and resized them into $224\times 224$. Note that, the randomly crop and flip operations are also adopted for data augmentation.
@@
During the training phase, we utilize the ground truth as the input and predict the attributes in a parallel manner, which can reduce the cumulative error significantly and make our model converge faster. For the model inference, we adopt a step-by-step generation approach that uses the output predicted in the previous step and historical output as the inputs of our network. For the detailed parameters, we set the learning rate as 1e-5. We train our model for 50 epochs with the Adam optimizer and set the batch size as 32. Our source code is implemented using Python and the deep learning framework PyTorch [21]. The experiments are conducted based on a server with GPU A100. More details can be found in our source code.
4.3 Comparison with Other SOTA Models
@@
In this section, we will report our recognition results on all six datasets and compare them with existing state-of-the-art pedestrian attribute recognition algorithms. Note that the results of VTB* are obtained by replacing the backbone network of VTB using the visual encoder of the CLIP model.
@@
Results on PETA dataset.  As shown in Table 1, our proposed SequencePAR model achieves the best performance on most of the evaluation metrics, i.e., 84.92, 90.44, 90.73, 90.46 on the Accuracy, Precision, Recall, and F1-measure, respectively. Compared with other Transformer-based pedestrian attribute recognition models, we can find that our model exceeds both the VTB (ViT-base//16, 79.60, 86.76, 87.17, 86.71) and VTB* (ViT-L//14, 79.59, 86.66, 87.82, 86.97) by +5.32//+3.68//+3.56//+3.67 and +5.33// +3.78//+2.91//+3.41. Our model also achieves better performance than the DRFormer [28] on the Accuracy, Precision, and F1-measure metric. Therefore, we can draw the conclusion that our model achieves state-of-the-art results on the PETA dataset, which fully validates the superiority of our sequence generation framework for the pedestrian attribute recognition task.
@@
Results on PA100K dataset.  As shown in Table 1, our model SequencePAR achieves 83.94, 90.38, 90.23, and 90.10 in Accuracy, Precision, Recall, and F1, respectively. These results surpass those of the multi-label classification model VTB* (81.76, 87.87, 90.67, and 88.86), which also uses the ViT-L//14 model as the backbone. These experiments demonstrate that our attribute generation-based approach outperforms the discriminative-based learning framework. Additionally, our model achieves superior results compared to the state-of-the-art recognition models.
@@
Results on RAPv1 dataset.  As shown in Table 2, our model achieves 71.47, 82.40, 82.09, and 82.02 in accuracy, precision, recall, and F1, respectively. In contrast, the strong baseline VTB*, a discriminative method, achieves 69.78, 78.09, 85.21, and 81.10 on these metrics. Our results clearly outperform those of the baseline and other compared attribute recognition models, including DAFL [11] (AAAI-2022) and DRFormer [28] (NC 2022). These experiments comprehensively validate the effectiveness of our proposed model for pedestrian attribute recognition.
@@
Results on RAPv2 dataset.  As shown in Table 2, the multi-label classification-based attribute recognition model VTB* (backbone: ViT-L//14) achieves 67.58, 76.19, 84.00, and 79.52 on the four metrics, respectively. In comparison, our model achieves 70.14, 81.37, 81.22, and 81.10. Therefore, we can conclude that our model outperforms the strong discriminative learning-based pedestrian attribute recognition algorithm.
@@
Results on PETA-ZS dataset.  In addition to the aforementioned standard evaluation, we also conduct experiments on zero-shot setting datasets in this work. According to the experimental results reported in Table 3, our model achieves 66.70, 78.75, 78.52, and 78.40 in Accuracy, Precision, Recall, and F1, respectively. These results outperform those of the multi-label classification model VTB* (63.12, 74.77, 77.24, 75.50), which also uses the ViT-L//14 model as the backbone network.
@@
Results on RAP-ZS dataset.  The experimental results reported on the RAP-ZS dataset, as illustrated in Table 3, show that the baseline method VTB* achieves 68.34, 76.81, 84.51, and 80.07 in Accuracy, Precision, Recall, and F1, respectively. However, our model achieves better results on these metrics, namely 70.28, 82.13, 80.55, and 81.14. Experiments on the PETA-ZS and RAP-ZS datasets also show that our model performs better in this setting. We attribute this to the use of a large visual-language model, which exhibits better generalization ability. Another key reason is that the generative learning-based strategy is more effective for pedestrian attribute recognition.
@@
Table 4: Compare the greedy search and beam search with different beam widths. The best results are highlighted in bold.
@@
4.4 Ablation Study
Greedy Search or Beam Search for Attribute Generation? For the sequence generation tasks, like machine translation and image//video captioning, the greedy search and beam search are all widely used in the inference phase. The greedy search policy will select and remain the element with the maximum response score. In contrast, the beam search policy will always select and retain a fixed number of candidates (also called beam width $\mathcal{B}$) for each time step, and select the trajectory with the maximum summarised response.
In this part, we evaluate different decoding strategies to verify their effectiveness for pedestrian attribute generation. As shown in Table 4, the greedy search achieves $84.74$, $90.38$, $90.42$, and $90.28$ on Accuracy, Precision, Recall, and F1, respectively. When beam search is adopted, the recognition performance on the PETA dataset remains comparable to that of greedy search, but no further improvement is observed. Therefore, we employ the simple greedy search strategy for all subsequent experiments in this work.
Results using Different Layers of Decoders. As shown in Table 5, we conduct experiments on the PETA dataset to examine the impact of varying the number of layers in our decoder network. Specifically, we set the number of layers to 1, 3, 6, 9, and 12. Table 5 shows that the F1 scores are 89.80, 89.68, 90.46, 89.60, and 89.76, demonstrating relative stability. The best recognition performance on the PETA dataset is achieved when six decoding layers are used. Therefore, we use six layers as the default configuration for our decoder network in subsequent experiments.
Table 5: Comparison of Different Decoding Layers on PETA dataset. The best results are highlighted in bold.